Daniel is currently 26 years older than his son. In six years he will be three times older than his son. How old are both of them now?
Answer by lwsshak3(11628)

Daniel is currently 26 years older than his son. In six years he will be three times older than his son. How old are both of them now?
***
let x=son's age
x+26=Daniel's age
..
In 6 yrs:
x+6=son's age
x+26+6=x+32
x+32=3(x+6)
x+32=3x+18
2x=14
x=7
x+26=33
son's age=7
Daniel's age=33